Return to question hours. At the beginning of the year, the company provided the following estimates: 3.35 points DepartmentReturn to ques Required: 1. Compute the predetermined overhead rates used in the Molding Department and the Painting DepartmeAnswer is not complete. Complete this question by entering your answers in the tabs below. Reg 1 Reg 2 Reg 3A Reg 3B ComputeComplete this question by entering your answers in the tabs below. Reg 1 Req 2 Reg 3A Reg 3B What would be the total manufact3-a. wnat would be the total manufacturing cost recordea for JOD 205! 3-b. If the job contained 35 units, what would be the u
Answer #1
Calculation of pre determined overhead rate
Molding Painting
Fixed manufacturing overhead cost 193,600 485,940
Direct labor hours 33,000 53,400
Machine hours 88,000 31,000
Fixed manufacturing overhead rate 2.2 9.1
Variable manufacturing overhead rate 3.2 5.2
Pre determined overhead rate 5.4 14.3
2.Overhead cost applied = 370*5.4 + 129*14.3 3842.7
3.Total manufacturing cost Molding Painting Total
Direct material 932.00 1220.00 2152.00
Direct Labor 700.00 970.00 1670.00
Overhead cost 1998.00 1844.70 3842.70
Total manufacturing cost 7664.70
3-B Unit product cost = 7664.70/35 = $218.99 i.e. $219